OpenStudy (carrionbeast):

Can Somebody please convert this equation to a logarithmic equation? 42.07e^(0.75x) = y Edit: "e" as in the constant e equal to 2.718218 and so on.

jhonyy9 (jhonyy9):

i think may be ,,e" the base of ligarithm naturale - yes ?

OpenStudy (carrionbeast):

I edited the question jhonyy9.

jhonyy9 (jhonyy9):

where or how ? i dont see there nothing modified ?

jhonyy9 (jhonyy9):

first of all divide both sides by 42,07 so what will get ?

OpenStudy (carrionbeast):

"e" as in the constant e equal to 2.718218 and so on. y/42.07 = e^-0.75x Does this work or do I need to multiply 42.07 by the constant e first?

OpenStudy (carrionbeast):

Then would it be Ln(y/42.07) = -0.75x? To convert that into a function that can be graphed using technology you use the change-base formula to make it Log(y/42.07)/Log(e)? I'm lost here.

jhonyy9 (jhonyy9):

but firtsly you wrote e^(0,75x) how get this minus ?

OpenStudy (carrionbeast):

My bad, that was a typo. The original equation is y = 42.07e^(-0.75x).

jhonyy9 (jhonyy9):

but you need to solve it for x ?

OpenStudy (carrionbeast):

I just need to make it a logarithmic equation.

OpenStudy (phi):

y = 42.07e^(-0.75x) take at the ln of both sides: ln y = ln(42.07) - 0.75 x we could write that as 0.75 x = - ln y + ln 42.07 or x= 4/3 ln(y) + 4/3 ln(42.07)

OpenStudy (carrionbeast):

Why does one take the ln of both sides?

OpenStudy (phi):

to keep it an equation, you have to "do the same thing" to both sides

OpenStudy (carrionbeast):

Does the ln get rid of the constant e? Why did the exponent become a term by itself?

OpenStudy (phi):

let me fix the typo: or x= -(4/3) ln(y) + (4/3) ln(42.07)

OpenStudy (phi):

by definition ln(e^x) = x also, there is this property log(ab) = log(a) + log(b) (for any base, which includes ln i.e. base e) so ln(4 e^x) = ln(4) + ln(e^x) = ln(4) + x

OpenStudy (carrionbeast):

What level of math did you learn this in because my Algebra 2 teacher failed to mention the "ln(e^x) = x" definition.

OpenStudy (phi):

Did they mention \[ \log_b b^x =x \] ?

OpenStudy (carrionbeast):

Not specifically but I do know that the exponent on be can become a coefficient and then l\[ \log_{b} b = 1\] . In the end x=x

OpenStudy (phi):

ok, so \[ \log_b b^x = x \log_b b = x \cdot 1= x \] ln means \( \log_e \) so we do the same thing for \[ \ln e^x = x \log_e e = x \cdot 1 = x \]

OpenStudy (phi):

but to me "log" means "grab the exponent" as long as we have the "correct" base

OpenStudy (carrionbeast):

That is understandable. Thank you for the hard work.